OpenStudy (anonymous):

write the equation of the line with the slope of -3 and through points (5, -8)

jimthompson5910 (jim_thompson5910):

The general linear equation in slope intercept form is y = mx+b m is the slope b is the y-intercept

jimthompson5910 (jim_thompson5910):

y = mx+b y = -3x+b ... plug in the slope of -3 -8 = -3*(5) + b ... plug in the given point (x,y) = (5,-8) now we solve for b -8 = -3*(5) + b -8 = -15 + b -8 + 15 = b 7 = b b = 7 ---------------------------- So we know that m = -3 and b = 7 Therefore, y = mx+b turns into y = -3x + 7